OpenStudy (anonymous):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

y--2=2(x-3)

OpenStudy (anonymous):

slope intercept: to find this you have to... -2=3(2)+b -2=6+b -6 -6 -8=b now: y=2x-8

OpenStudy (anonymous):

Now standard: subtract the 2x -2x+y=-8

OpenStudy (anonymous):

Perfect! Thank you!

OpenStudy (anonymous):

No problem! Did you understand how I got that all?